OpenStudy (anonymous):

prove why f(x)=1/x uniformly continuous on (2,infinity )

OpenStudy (zzr0ck3r):

|f(x)-f(y)|=|1/x-1/y|=|(x-y)/(xy)|<|x-y|/a^2 let delta = a^2*epsilon then |f(x)-f(y)|<|x-y|/a^2<epsilon

OpenStudy (zzr0ck3r):

you could actually get rid of the a variable since x,y are greater than 1. |x-y|/xy < |x-y| then choose delta=epsilon and you are good....